# On the dynamics of leapfrogging vortex rings

### v1

@@ 6.2. Strong localization in the vertical direction
$\mathcal{U}_{Z}^{i,\varepsilon}=\{(r,z)\in{\mathbb{R}}^{2}_{+},\;|z-z^{*}|<Z\}$
@@
$\mu_{t}^{i,\varepsilon}(\eta,Z)=\int_{{\mathbb{R}}^{2}}\big{(}1-W_{Z,\eta}(x_{z}-z^{*})\Big{)}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x.$
@@
There exists $C>0$ independent of $T_{0}$ such that, for every $t\leqslant T_{\varepsilon}$, $\varepsilon>0$, $i\in\{1,\ldots,N\}$, $\eta>0$ and $Z\in\left[\frac{h_{0}}{128},\frac{h_{0}}{16}\right]$, we have that
@@
by reproducing the arguments of Lemma 6.3 with the only difference that $Z/2\geqslant\frac{h_{0}}{256}$, and
@@
provided $Z\in\left[\frac{h_{0}}{64},\frac{h_{0}}{16}\right]$. Injecting this into Lemma 6.8, we get that
@@
We take now $Z_{0}=\frac{d_{0}}{16}$ $\eta=\frac{Z_{0}}{2n}$ with $n=[|\ln\varepsilon|]-1$, and get that
@@
for all $Z\in\left[\frac{h_{0}}{64},\frac{h_{0}}{16}\right]$. Then by the same iterative method than in Lemma 6.4, we obtain that
$\mu_{t}^{\varepsilon}(Z_{0},\eta)\leqslant\frac{(Cnt)^{n}}{n!}.$
@@
$\varepsilon^{-\ell}\mu_{t}^{\varepsilon}(Z_{0},\eta)$ $\leqslant\frac{1}{\sqrt{\pi n}}\exp\Big{(}n\big{(}\ln(Cnt)-\ln(n)+1\big{)}+\ell|\ln\varepsilon|\Big{)}$
$\leqslant\frac{2}{\sqrt{\pi|\ln\varepsilon|}}\exp\Big{(}|\ln\varepsilon|\big{(}\ln(Ct)+1+\ell\big{)}\Big{)}$
which gives the result as soon as $CT_{0,\ell}\leqslant 1/e^{1+\ell}$. ∎
@@
By Lemmas 6.6 and 6.10, we have obtained the proof of Proposition 6.1 which implies to Part (ii) of Theorem 1.1, see (3.4) for the definition of $\mathcal{A}_{\varepsilon}^{r_{0}}(d,h_{0})$.
Proposition 6.1 also implies that $T_{\varepsilon}=T_{0}$, so all the estimates we obtained throughout this article hold on the time interval $[0,T_{0}]$ which is independent of $\varepsilon$. In particular Corollary 5.5 gives Part (i) of Theorem 1.1.
@@ Appendix A Two classical lemmas
First, we recall a standard lemma for mass rearrangement, whose proof is available in [31, Lemma B.1]. Let $g$ be a non increasing continuous function from $(0,+\infty)$, non-negative, such that $s\mapsto sg(s)\in L^{1}_{\mathrm{loc}}\big{(}[0,\infty)\big{)}$. Let
@@
The second lemma concerns a variant of Gronwall’s inequality, whose proof is provided in [20, Lemma A.2].
@@
The proof was already done in [20] and is recalled here for the sake of completeness.
@@ Appendix B On the suitable scaling regime
The purpose of this appendix is to provide a more detailed discussion of suitable scaling regime leading to the leapfrogging dynamics chosen in this paper as well as a detailed comparison to the results of [11]. While the authors obtain a result somewhat in the spirit of Theorem 1.1, the considered asymptotic regime differs significantly for the present one as does the limiting trajectory.
@@ B.1. The scaling regime and respective limiting trajectory
hence depending on the reference radius $r^{\ast}$, the intensity $\gamma$ and in logarithmically diverging manner on its thickness $\varepsilon$. When several vortex rings are considered, their motion is driven by the self-induced motion and their mutual interactions which are known to be of strength that is inversely proportional to their mutual distance. Depending on the chosen asymptotic regime, the former [12, 8] or the latter [10] maybe dominant. The leapfrogging regime corresponds to a regime balancing these two effects being critical in a sense and the rigorous justification of the asymptotic dynamics is significantly more difficult. Indeed, for the 3D Euler equations it remained completely open until the recent aforementioned [18, 11] and our work. As a rule of thumb and outlined in the introduction, the leapfrogging dynamics is expected to occur for vortex rings that travel at leading order at the same self-induced velocity (1.4) of order $|\ln\varepsilon|$, namely share the same reference radius $r^{\ast}$ and intensity $\gamma$. Further, their mutual distance is of order $1/{\sqrt{|\ln\varepsilon|}}$ leading to a mutual interaction of order $\sqrt{|\ln\varepsilon|}$. In a time of order $\mathcal{O}\left(1/|\ln\varepsilon|\right)$ (so of order 1 in the rescaled time variable), this, in turn, generates a variation of the rings’ radii of order $\mathcal{O}\left(1/{\sqrt{|\ln\varepsilon|}}\right)$ which implies a change in their self-induced velocity (1.4) at the order $\sqrt{|\ln\varepsilon|}$, i.e.
$v_{L}=|\ln\varepsilon|\frac{\gamma}{4\pi r^{*}}e_{z}+\mathcal{O}(\sqrt{|\ln\varepsilon|})$
@@
Moreover, formally assuming that (1.9), namely $X_{i}^{\varepsilon}(t)=X^{*}+\frac{\gamma}{4\pi r^{*}}te_{z}+\frac{1}{\sqrt{|\ln\varepsilon|}}Y_{i}^{\varepsilon}(t)$, holds, we derive the motion law for $Y_{i}^{\varepsilon}(t)$ with $Y_{i}^{\varepsilon}(0)=Y_{i,0}$ obtaining that
@@
$=\frac{\gamma}{2\pi}\sum_{j\neq i}\frac{(Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon})^{\perp}}{|Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon}|^{2}}-\frac{\gamma}{4\pi(r^{\ast})^{2}}e_{z}+\mathcal{O}\left(\frac{1}{\sqrt{|\ln\varepsilon|}}\right).$
We recognize that the relative self-induced motion and interaction are of the same order. We refer to Section 5.2 for full details. In particular, it follows from Lemma 5.3 that the limit trajectories $\widetilde{Y}_{i}$ defined in (1.9) satisfy (1.10), i.e.
@@
We refer to [18, (1.14)] and [33, Section 1.2] and references therein for a qualitative and quantitative analysis of solutions to (B.1) and prove in Appendix C that no collisions occur in finite time.
@@ B.2. Comparison of scaling regimes in the literature
Different from the present the one, see also [18, 35], the authors in [11] consider a scaling regime in which the self-induced motion of the vortex rings is of the same order as the velocity induced by the interaction of vortex rings. We also refer the reader to [42, Section 1.1.1-1.1.2.] for a comparison of both regimes. More precisely, the authors in [11] work with vortex rings of radius $\mathcal{O}(|\ln\varepsilon|)$, vortex circulations of order $\mathcal{O}(1)$ and the vorticity $\Omega$ that reads in terms of the potential vorticity $\omega=\frac{\Omega}{r}$, as follows
$\begin{cases}\partial_{t}\frac{\Omega}{r}+u\cdot\nabla\left(\frac{\Omega}{r}\right)=0,\\ u=\frac{1}{r}\nabla^{\perp}\Psi,\\ \nabla\cdot\left(\frac{1}{r}\nabla\Psi\right)=\Omega\\ \int\Omega_{i}\mathrm{d}x=a_{i},\\ \operatorname{supp}\Omega_{i}\subset B((0,\alpha|\ln\varepsilon|)+x_{i}),\end{cases}$
$\begin{cases}\partial_{t}\omega+u\cdot\nabla\omega=0,\\ u=\frac{1}{r}\nabla^{\perp}\Psi,\\ \nabla\cdot\left(\frac{1}{r}\nabla\Psi\right)=r\omega\\ \int r\omega_{i}\mathrm{d}x=a_{i},\\ \operatorname{supp}\omega_{i}\subset B((0,\alpha|\ln\varepsilon|)+x_{i}).\end{cases}$
@@
and check that the vortex rings’ circulation is order $\mathcal{O}(1)$, specifically
@@
so that the scaled rings’ radius of order $\mathcal{O}(1)$. The distances between the rings are now of order $\mathcal{O}(\frac{1}{|\ln\varepsilon|})$, which is more singular than in (1.6) of our analysis. We then compute that
@@
We then turn to the Biot-Savart law. Our goal is to determine $\widetilde{\Psi}$ such that
$\nabla\cdot\left(\frac{1}{r}\nabla\widetilde{\Psi}\right)=r\widetilde{\omega}.$
@@
$\nabla\cdot\left(\frac{1}{r}\nabla\widetilde{\Psi}\right)(t,x)$ $=\frac{1}{|\ln\varepsilon|}\left(-\frac{|\ln\varepsilon|e_{r}}{r^{2}}\cdot\nabla\Psi(t,|\ln\varepsilon|x)+\frac{|\ln\varepsilon|^{2}}{r}\Delta\Psi(t,|\ln\varepsilon|x)\right)$
$=\frac{|\ln\varepsilon|^{3}}{|\ln\varepsilon|}\nabla\cdot\left(\frac{1}{r}\nabla\Psi\right)(t,|\ln\varepsilon|x)$
@@
In conclusion, the results of [11] can be interpreted as strong confinement result (in both directions) for vortex rings of radius of order $\mathcal{O}(1)$ and circulations of order $\mathcal{O}(1)$, placed at a distance of order $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$, for a time of order $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|^{2}}\right)$. Indeed, one can perform a final change of the time variable
@@
Finally, the scaled potential vorticity satisfies the transport equation
@@
This regime corresponds to the second regime in [42, Section 1.1.2.] where it is expected that the two rings have the same local velocities. The velocities $v_{L}$ (local induction approximation) and $F$ (generated by the other rings) are both of order $|\ln\varepsilon|$. This means that in this regime the vortex rings perform a rotation around each other over a distance of order $\mathcal{O}\left(\frac{1}{|\ln\varepsilon}\right)$ in a time of order $1/|\ln\varepsilon|^{2}$. Moreover, in that regime, the vortex rings can have different circulations $\gamma_{i}$, and in that case the limiting trajectories were proven [11, Theorem 2.1] to satisfy
@@
Note that (B.4) differs from (B.1) in terms of the vertical translation. The properties of solutions to (B.4) have been investigated in [39], in particular ruling out blow-up provided that all $\gamma_{i}$ are of the same sign [39, Theorem 2.1.]. See also [11, Section 7] for the case of two vortex rings.
@@
our scaling $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(|\ln\varepsilon|)$ $\mathcal{O}\left(\frac{1}{\sqrt{|\ln\varepsilon|}}\right)$ $\mathcal{O}(\sqrt{|\ln\varepsilon|})$ $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$
@@
Appendix C No collision in finite time
In this section we prove that no collisions in finite time may occur for system (1.10), rewritten here as

### v2

@@ 6.2. Strong localization in the vertical direction
$\mathcal{U}_{Z}^{i,\varepsilon}=\{(z,r)\in{\mathbb{R}}^{2}_{+},\;|z-z^{*}|<Z\}$
@@
$\mu_{t}^{i,\varepsilon}(Z,\eta)=\int_{{\mathbb{R}}^{2}}\big{(}1-W_{Z,\eta}(x_{z}-z^{*})\Big{)}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x.$
@@
There exists $C>0$ independent of $T_{0}$ such that, for every $t\leqslant T_{\varepsilon}$, $\varepsilon>0$, $i\in\{1,\ldots,N\}$, $\eta>0$ and $Z\in\left[\frac{h_{0}}{32},\frac{h_{0}}{16}\right]$, we have that
@@
by reproducing the arguments of Lemma 6.3 with the only difference that $Z/2\geqslant\frac{h_{0}}{64}$, and
@@
provided $Z\in\left[\frac{h_{0}}{32},\frac{h_{0}}{16}\right]$. Injecting this into Lemma 6.8, we get that
@@
We take now $Z_{0}=\frac{d_{0}}{16}$ $\eta=\frac{Z_{0}}{2n}$ with $n=[|\ln\varepsilon|]$, and get that
@@
for all $Z\in\left[\frac{h_{0}}{32},\frac{h_{0}}{16}\right]$. Then by the same iterative method than in Lemma 6.4, we obtain that
$\mu_{t}^{\varepsilon}(Z_{0}-\eta,\eta)\leqslant\frac{(Cnt)^{n-1}}{(n-1)!}.$
@@
$\varepsilon^{-\ell}\mu_{t}^{\varepsilon}(Z_{0}-\eta,\eta)$ $\leqslant\frac{1}{\sqrt{\pi(n-1)}}\exp\Big{(}(n-1)\big{(}\ln(Cnt)-\ln(n-1)+1\big{)}+\ell|\ln\varepsilon|\Big{)}$
$\leqslant\frac{2}{\sqrt{\pi|\ln\varepsilon|}}\exp\Big{(}|\ln\varepsilon|\big{(}\ln(Ct)+2+\ell\big{)}\Big{)}$
which gives the result as soon as $CT_{0,\ell}\leqslant 1/e^{2+\ell}$. ∎
@@
By proving Lemmas 6.6 and 6.10, we have completed the proof of Proposition 6.1 which implies Part (ii) of Theorem 1.1, see (3.4) for the definition of $\mathcal{A}_{\varepsilon}^{r_{0}}(d,h_{0})$.
Proposition 6.1 also implies that $T_{\varepsilon}=T_{0}$, so all the estimates we obtained throughout this article hold on the time interval $[0,T_{0}]$ which is independent of $\varepsilon$. In particular, we recall that Corollary 5.5 yields Part (i) of Theorem 1.1.
@@ Appendix A Two classical lemmas
First, we recall a standard lemma for mass rearrangement, whose proof is available in [35, Lemma B.1]. Let $g$ be a non increasing continuous function from $(0,+\infty)$, non-negative, such that $s\mapsto sg(s)\in L^{1}_{\mathrm{loc}}\big{(}[0,\infty)\big{)}$. Let
@@
The second lemma concerns a variant of Gronwall’s inequality, whose proof is provided in [23, Lemma A.2].
@@
The proof was already provided in [23, Lemma A.2] and is recalled here for the sake of completeness.
@@ Appendix B On the suitable scaling regime
The purpose of this appendix is to provide a more detailed discussion on the suitable scaling regime leading to the leapfrogging dynamics chosen in this paper as well as a detailed comparison to the results of [11]. While the authors obtain a result somewhat in the spirit of Theorem 1.1, the considered asymptotic regime differs significantly for the present one as does the limiting trajectory.
@@ B.1. The scaling regime and respective limiting trajectory
hence depending on the reference radius $r^{\ast}$, the intensity $\gamma$ and in logarithmically diverging manner on its thickness $\varepsilon$. When several vortex rings are considered, their motion is driven by the self-induced motion and their mutual interactions which are known to be of strength that is inversely proportional to their mutual distance. Depending on the chosen asymptotic regime, the former [12, 8] or the latter [10] may be dominant. The leapfrogging regime corresponds to a regime balancing these two effects and the rigorous justification of the asymptotic dynamics is significantly more difficult. Indeed, for the 3D Euler equations it remained completely open until the recent aforementioned [20, 11] and our work. As a rule of thumb and outlined in the introduction, the leapfrogging dynamics is expected to occur for vortex rings that travel at leading order at the same self-induced velocity (1.1) of order $|\ln\varepsilon|$, namely share the same reference radius $r^{\ast}$ and intensity $\gamma$. Further, their mutual distance is of order $1/{\sqrt{|\ln\varepsilon|}}$ leading to a mutual interaction of order $\sqrt{|\ln\varepsilon|}$. In a time of order $\mathcal{O}\left(1/|\ln\varepsilon|\right)$ (so of order 1 in the rescaled time variable), this, in turn, generates a variation of the rings’ radii of order $\mathcal{O}\left(1/{\sqrt{|\ln\varepsilon|}}\right)$ which implies a change in their self-induced velocity (1.1) at the order $\sqrt{|\ln\varepsilon|}$, i.e.
$v_{L}=|\ln\varepsilon|\frac{\gamma}{4\pi r^{*}}e_{z}+\mathcal{O}(\sqrt{|\ln\varepsilon|}).$
@@
Moreover, defining the variation in the frame moving at constant speed, namely setting $X_{i}^{\varepsilon}(t)=X^{*}+\frac{\gamma}{4\pi r^{*}}te_{z}+\frac{1}{\sqrt{|\ln\varepsilon|}}Y_{i}^{\varepsilon}(t)$, we derived in Lemma 5.3 the motion law for $Y_{i}^{\varepsilon}(t)$ with $Y_{i}^{\varepsilon}(0)=Y_{i,0}$:
@@
$=\frac{\gamma}{2\pi}\sum_{j\neq i}\frac{(Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon})^{\perp}}{|Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon}|^{2}}-\frac{\gamma}{4\pi(r^{\ast})^{2}}Y_{i,r}^{\varepsilon}e_{z}+\mathcal{O}\left(\frac{1}{\sqrt{|\ln\varepsilon|}}\right).$
We recognize that the relative self-induced motion and interaction are of the same order. We refer to Section 5.2 for full details. In particular, we recover in this regime that smallest rings goes faster (see also (1.1) for this well-known fact).
Note that as $\varepsilon\rightarrow 0$, see also Lemma 5.3, one recovers the following system for the limiting trajectories
@@
We refer to [20, Equation (1.14)] and [39, Section 1.2] and references therein for a qualitative and quantitative analysis of solutions to (B.1). Further, we prove in Appendix C.1 that no collisions occur in finite time for (B.1). Appendix C.2 elucidates differences in the behavior of solutions $\widetilde{Y}_{i}$ to (B.1) ($\varepsilon=0$) and $Y_{i}^{\varepsilon}$ as defined in (1.3) to (1.8) ($\varepsilon>0$).
@@ B.2. Comparison of scaling regimes in the literature
Different from the present the one, see also [20, 39], the authors in [11] consider a scaling regime in which the self-induced motion of the vortex rings is of the same order as the velocity induced by the interaction of vortex rings. We also refer the reader to [48, Section 1.1.1-1.1.2] for a comparison of both regimes. More precisely, the authors in [11] work with vortex rings of radius $\mathcal{O}(|\ln\varepsilon|)$, vortex circulations of order $\mathcal{O}(1)$ and the vorticity $\Omega$ that reads in terms of the potential vorticity $\omega=\frac{\Omega}{r}$, as follows
$\begin{cases}\partial_{t}\frac{\Omega}{r}+u\cdot\nabla\left(\frac{\Omega}{r}\right)=0,\\ u=\frac{1}{r}\nabla^{\perp}\Psi,\\ \operatorname{div}\left(\frac{1}{r}\nabla\Psi\right)=\Omega\\ \int\Omega_{i}\mathrm{d}x=a_{i},\\ \operatorname{supp}\Omega_{i}\subset B((0,\alpha|\ln\varepsilon|)+x_{i}),\end{cases}$
$\begin{cases}\partial_{t}\omega+u\cdot\nabla\omega=0,\\ u=\frac{1}{r}\nabla^{\perp}\Psi,\\ \operatorname{div}\left(\frac{1}{r}\nabla\Psi\right)=r\omega\\ \int r\omega_{i}\mathrm{d}x=a_{i},\\ \operatorname{supp}\omega_{i}\subset B((0,\alpha|\ln\varepsilon|)+x_{i}).\end{cases}$
@@
and check that the vortex rings’ circulation is of order $\mathcal{O}(1)$, specifically
@@
so that the scaled rings’ radii are of order $\mathcal{O}(1)$. The distances between the rings are now of order $\mathcal{O}(\frac{1}{|\ln\varepsilon|})$, which is more singular than in (1.4) of our analysis. We then compute that
@@
Turning to the Biot-Savart law, we seek to determine $\widetilde{\Psi}$ such that
$\operatorname{div}\left(\frac{1}{r}\nabla\widetilde{\Psi}\right)=r\widetilde{\omega}.$
@@
$\operatorname{div}\left(\frac{1}{r}\nabla\widetilde{\Psi}\right)(t,x)$ $=\frac{1}{|\ln\varepsilon|}\left(-\frac{|\ln\varepsilon|e_{r}}{r^{2}}\cdot\nabla\Psi(t,|\ln\varepsilon|x)+\frac{|\ln\varepsilon|^{2}}{r}\Delta\Psi(t,|\ln\varepsilon|x)\right)$
$=\frac{|\ln\varepsilon|^{3}}{|\ln\varepsilon|}\operatorname{div}\left(\frac{1}{r}\nabla\Psi\right)(t,|\ln\varepsilon|x)$
@@
In conclusion, the results of [11] can be interpreted as a strong confinement result (in both directions) for vortex rings of radius of order $\mathcal{O}(1)$ and circulations of order $\mathcal{O}(1)$, placed at a distance of order $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$, for a time of order $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|^{2}}\right)$. Indeed, one can perform a final change of the time variable
@@
to recover the transport equation
@@
This regime corresponds to the second regime in [48, Section 1.1.2] where it is expected that the two rings have the same local velocities. The velocities $v_{L}$ (local induction approximation) and $F$ (generated by the other rings) are both of order $|\ln\varepsilon|$. This means that in this regime the vortex rings perform a rotation around each other over a distance of order $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$ in a time of order $1/|\ln\varepsilon|^{2}$. Moreover, in that regime, the vortex rings can have different circulations $\gamma_{i}$, and in that case the limiting trajectories were proven [11, Theorem 2.1] to satisfy
@@
Note that (B.4) differs from (B.1) in terms of the vertical translation, and we remark that two vortex rings with the same circulation have the same local induction velocity, no matter of their radius. The properties of solutions to (B.4) have been investigated in [43], in particular ruling out blow-up provided that all $\gamma_{i}$ are of the same sign [43, Theorem 2.1]. See also [11, Section 7] for the case of two vortex rings.
@@
our scaling $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\frac{|\ln\varepsilon|\gamma}{4\pi r^{*}}e_{z}+\mathcal{O}(\sqrt{|\ln\varepsilon|})$ $\mathcal{O}\left(\frac{1}{\sqrt{|\ln\varepsilon|}}\right)$ $\mathcal{O}(\sqrt{|\ln\varepsilon|})$ $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$
@@
Because of their mutual distances and the intensity of their velocities, it should be noted that both time scales are respectively suitable for observing a rotation of one vortex ring with respect to the other.
Appendix C Some comments on the limit dynamics
As detailed in the introduction, we have considered two limiting trajectories: $(Y_{i}^{\varepsilon})_{i=1,\dots,N}$ defined in (1.7) and $(\widetilde{Y}^{i})_{i=1,\dots,N}$ obeying (1.8), and we have proved in Lemma 5.3 that they remain close, namely at distance $\mathcal{O}(1/\sqrt{|\ln\varepsilon|})$. We first remark that there is no-collision in finite time for $(\widetilde{Y}^{i})$ and that they are global in time. Next, we present some numerical evidence on the difference of these two limiting dynamics. Let us mention that the dynamics $(Y_{i}^{\varepsilon})$ also differs from the dynamics studied in [39], even though the two dynamics have the same limiting trajectories $(\widetilde{Y}_{i}(t))$.
C.1. No collision in finite time and global solutions for the second limiting system
For simplicity, we rewrite System (1.8) here as